Solve the following system by substitution
5x-2y=-5
y-5x=3

Answer by jainenderkapoor(61)

5x-2y=-5 ----(1)
y-5x=3 ----(2)
From the second equation
we have
y - 5x = 3
y - 5x + 5x = 3 + 5x (adding 5x to both sides)
y = 3 + 5x ---- (3)
substituting this value in (1) equation, we get
5x - 2(3 + 5x) = -5
5x - 6 - 10x = -5
-5x - 6 = -5
-5x -6 + 6 = -5 +6 (adding 6 to both sides)
-5x = 1
-5x/-5 = 1/-5 (dividing both sides by -5)
x = -1/5
Putting this value of x in (3) equation we get
y = 3 + 5(-1/5)
y = 3 - 1
y = 2
so solution is x = -1/5 and y = 2
Hope the solution is clear to you. In case you have any other doubt, please contact me on jainenderkapoor@gmail.com I am available for Online Tutoring at very reasonable rates.